# Finance as Extended Biology: Reciprocity as the Cognitive Substrate of Financial Behavior

### v1

@@ Abstract
A central challenge in economics and artificial intelligence is explaining how financial behaviors—such as credit, insurance, and trade—emerge from decentralized social interaction. We argue that these functions are not institutionally engineered, but arise from a single behavioral substrate: reciprocity. Observed across primates and foundational to early human societies, reciprocity governed the exchange of goods, regulation of obligation, and maintenance of cooperation—without formal rules. Trade, widely regarded as the origin of financial structure, is reframed here as the canonical form of reciprocity: simultaneous, symmetric, and partner-contingent. Building on this logic, we reconstruct four fundamental financial functions—credit, insurance, token exchange, and investment—as structured extensions of the same interaction principle. By grounding financial behavior in minimal, simulateable social dynamics, this framework shifts the explanatory focus from institutional design to behavioral foundations, offering a new basis for modeling decentralized cooperation in both human and artificial systems.
@@ 3.3 Toward a Behavioral Substrate: From Correction to Construction
We propose reciprocity as such a substrate. Reciprocal exchange is not exclusive to humans: among non-human primates, it reliably emerges across social contexts, even without formal rules or enforcement mechanisms De Waal (1997). In early human societies, however, reciprocity assumed a far more central role, becoming the organizing principle of economic life—governing the circulation of goods, the distribution of obligations, and the maintenance of long-term relationships. Long before markets, money, or states, reciprocity structured how communities coordinated cooperation, regulated access, and sustained social cohesion—not as an abstract ideal, but as an empirically documented foundation of real-world exchange Mauss (2024); Sahlins (2013); Malinowski (2013).
@@ 5 A Simulateable Architecture for the Emergence of Financial Behavior
Building on prior work modeling simulateable reciprocity Diau (2025), a lightweight agent architecture is introduced for simulating the emergence of financial behavior from bottom-up interactions. Rather than relying on engineered incentives or explicit cooperation policies, the architecture specifies a minimal behavioral substrate—centered on partner recognition, reciprocal credence, and cost–return evaluation.
@@
Detecting Emergent Structure.
To evaluate whether financial behavior has emerged, we define system-level markers corresponding to the four core functions. These indicators are derived from observable interaction patterns and can be quantified through agent logs:

### v2

@@ Abstract
A central challenge in economics and artificial intelligence is explaining how financial behaviors—such as credit, insurance, and trade—emerge without formal institutions. We argue that these functions are not products of institutional design, but structured extensions of a single behavioral substrate: reciprocity. Far from being a derived strategy, reciprocity served as the foundational logic of early human societies—governing the circulation of goods, regulation of obligation, and maintenance of long-term cooperation well before markets, money, or formal rules. Trade, commonly regarded as the origin of financial systems, is reframed here as the canonical form of reciprocity: simultaneous, symmetric, and partner-contingent. Building on this logic, we reconstruct four core financial functions—credit, insurance, token exchange, and investment—as expressions of the same underlying principle under varying conditions. By grounding financial behavior in minimal, simulateable dynamics of reciprocal interaction, this framework shifts the focus from institutional engineering to behavioral computation—offering a new foundation for modeling decentralized financial behavior in both human and artificial agents.
@@ 3.3 Toward a Behavioral Substrate: From Correction to Construction
We identify reciprocity as the minimal substrate for decentralized financial behavior. It is not unique to humans: among non-human primates, reciprocal exchange emerges reliably across social contexts, even without formal rules or enforcement De Waal (1997). In early human societies, reciprocity became more than a behavioral tendency—it served as the organizing logic of economic life. It governed how communities exchanged goods, distributed obligations, and maintained long-term coordination. Long before markets or money, reciprocity structured economic relationships—not as an idealized norm, but as a documented feature of real-world exchange Mauss (2024); Sahlins (2013); Malinowski (2013).
@@
Financial Function Reciprocity Extension Behavioral Logic of Finance
Credit Delayed reciprocation I help you today, and you return the favor later.
Insurance Mutual support under uncertainty I help you now, trusting you’ll help me if I’m ever in trouble.
Token Exchange Indirect reciprocation via tokens I help a stranger, and someone else helps me later—with a token to keep track.
Investment Future-oriented reciprocation I give you something now, expecting you’ll return more in the future.
Table 1: Each financial function reflects a different way of extending reciprocity—adjusting how we give and receive across time, risk, and social context.
@@ 5 A Simulateable Architecture for the Emergence of Financial Behavior
Building on prior work modeling simulateable reciprocity Diau (2025), we propose a lightweight yet extensible architecture for simulating the emergence of financial behavior in multi-agent systems.
While the core behavioral substrate remains grounded in three primitives—partner recognition, reciprocal credence, and cost–return sensitivity—these mechanisms can be implemented in LLM-based agents via structured memory and simple interaction protocols, rather than reinforcement learning. This allows agents to reason about reciprocity and adjust their behavior without engineered rewards or fixed roles.
@@
Evaluating Simulation Results via Grounded Behavioral Criteria.
In multi-agent simulations—especially those involving LLM-based agents—there is often no clear ground truth, and existing evaluations rely on synthetic tasks or arbitrary metrics that lack correspondence to real-world behavior.
We address this by proposing a set of core behavioral criteria—credit, insurance, token exchange, and investment—that reflect well-documented dynamics in actual economic systems. These grounded, interpretable patterns are derived from observable interaction logs and serve as criteria for assessing whether meaningful financial structures have emerged: